I am trying to solve
-9(x)-3)^2=-7
Thanks

Answer by tutorcecilia(2144)

-9(x)-3^2=-7
-9x-9=-7[Use the order of operations (PEMDAS)]
-9x-9+9=-7+9 [solve for x]
-9x=2
x=-2/9
.
Check by plugging (x=-2/9) back into the original equation.